OpenStudy (firekat97):

How would you integrate x.e^(6x) by parts? :/

hartnn (hartnn):

have you tried integrating by parts before? What have you tried for this question? :)

OpenStudy (firekat97):

It is actually my first "real" attempt lol. So what I did so far, was to let x = u and e^(6x).dx = dv then I got du = dx and v = e^(6x)/6 After that I tried subbing everything back into uv - ∫v. du and I think this is where I messed up, but I'm not too sure actually :S

hartnn (hartnn):

your steps are correct, let subb what you got \(\large x \dfrac{e^{6x}}{6} - \int \dfrac{e^{6x}}{6} dx \)

hartnn (hartnn):

good so far?

OpenStudy (firekat97):

Omg! Wait nvm, I just solved it :D silly mistake lol, but thanks so much!! :D

hartnn (hartnn):

welcome ^_^ and since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (firekat97):

hahahaha ty :)